# WiMANS: A Benchmark Dataset for WiFi-based Multi-user Activity Sensing

### 24 Jan 2024 v1

@@ E.7 Maintenance
  7. If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to dataset consumers? If so, please provide a description. Yes. Contributions to the dataset and benchmarks are highly appreciated. Researchers can extend the dataset by annotating videos with human body joints for WiFi-based pose estimation. Annotations can be added to the dataset by pulling requests on the repository or though emails to dataset managers. All contributions will be properly acknowledged.

### 12 Mar 2024 v2

@@ 0.E.7 Maintenance
  7. If others want to extend/augment/build on/contribute to the data-set, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to dataset consumers? If so, please provide a description. Yes. Contributions to the dataset and benchmarks are highly appreciated. Researchers can extend the dataset by annotating videos with human body joints for WiFi-based pose estimation. Annotations can be added to the dataset by pulling requests on the repository or though emails to dataset managers. All contributions will be properly acknowledged.